Fermi National Laboratory
 

Operation Summary for March 17, 2010 - March 19, 2010

 

Wednesday March 17

The day shift began with Operations monitoring the Tevatron (TeV) store 7680 (18.8E30) and the Recycler stash (403.4E10), with the Antiproton Source stacking (19.3E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the NuMI and MiniBooNE experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams.

During the day shift, the shift began with TeV personnel conducting crystal collimator studies. A barrier fence was repaired near SSB. RF personnel swapped out the modulator for a TeV RF station (TRF1) after the TeV studies. There was trouble with a TeV RF anode supply between 10:30 AM and 1 PM; Operations put the TeV through a dry squeeze during this time. Interlock technicians repaired the Meson MTest gate at MT6-2. Operations put the TeV into HEP shot setup and established store 7682 at 2:40 PM with an Average Initial Luminosity (AIL) of 322.64E30. Operations established beam to MTest at 2:50 PM.

During the evening shift, Booster personnel replaced a bad power supply for the Noker at 6 PM. Operations disabled the MTest beam at 6 PM. Operations reported a misfiring Booster extraction kicker (B:MKS04); Booster personnel turned it off and compensated for its loss with the other kickers.

Thursday March 18

The midnight shift began with Operations monitoring store 7682 (105.15E30) and the stash (339.4E10), with Pbar stacking (10.4E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, around 4 AM, B:MKS05 tripped off just before store 7682 was scheduled for termination. Booster personnel came in and repaired both B:MKS04 and B:MKS05. The MINOS service building lost power at 6:55 AM; operators and the Duty Electrician responded; the yard breaker and a ground fault circuit on a transformer tripped. The Duty Electrician had their emergency generator on by 7:49 AM.

During the day shift, store 7682 collided until 8:30 AM. CDF and D-Zero made collision hall accesses until 11:15 AM. Operations put the TeV into HEP shot setup and established store 7685 at 2:15 PM with an Average Initial Luminosity (AIL) of 376E30 (a new record). NuMI beam was re-established at 10:15 AM. MKS04 sparking issues were investigated and BRF12 modulator developed a problem requiring disconnection from the anode supply for repairs. During the evening shift, store 7685 collided all shift. BRF12 was repaired and tied back into the West Anode supply by 6:39 PM. Operations disabled the MTest beam at 6 PM. At 11:27 PM, an I- ion gauge tripped off and closed the I- vacuum valve. Beam returned by 11:45 PM.

Friday March 19

The midnight shift began with Operations monitoring store 7685 (109.05E30) and the stash (303.9E10), with Pbar stacking (11.6E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, store 7685 collided all shift. Beam was established to M-Test at 4:00 AM. Linac RF station 1 tripped hard at 5:30 AM and remained off through the end of shift; experts were called in to investigate.

The Plan for Friday and the Weekend

Repair LRF1, put in a new store, then stack, stash, and store throughout the weekend.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 03/19/2010    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory